# The VLLM Safety Paradox: Dual Ease in Jailbreak Attack and Defense

### v1

@@ Abstract
The vulnerability of Vision Large Language Models (VLLMs) to jailbreak attacks appears as no surprise. However, recent defense mechanisms against these attacks have reached near-saturation performance on benchmarks, often with minimal effort. This simultaneous high performance in both attack and defense presents a perplexing paradox. Resolving it is critical for advancing the development of trustworthy models. To address this research gap, we first investigate why VLLMs are prone to these attacks. We then make a key observation: existing defense mechanisms suffer from an over-prudence problem, resulting in unexpected abstention even in the presence of benign inputs. Additionally, we find that the two representative evaluation methods for jailbreak often exhibit chance agreement. This limitation makes it potentially misleading when evaluating attack strategies or defense mechanisms. Beyond these empirical observations, our another contribution in this work is to repurpose the guardrails of LLMs on the shelf, as an effective alternative detector prior to VLLM response. We believe these findings offer useful insights to rethink the foundational development of VLLM safety with respect to benchmark datasets, evaluation methods, and defense strategies.
@@ 1 Introduction
Jailbreak attacks, often referred to as red-teaming [41], serve as the most common method for assessing the harmlessness of LLMs and Vision-LLMs (VLLMs) [53, 32, 7]. They are designed to circumvent the built-in restrictions or safeguards within models [23], eliciting them to produce malicious outputs, such as content related to illegal activities, hate speech, and pornography.
@@
Compared to their LLM counterparts, the vulnerability of VLLMs to jailbreak attacks has garnered considerable attention until very recently [61, 38, 30]. Some initial methods inject high-risk content into images through typography or generative techniques like stable diffusion [45]. Leveraging such methods, datasets have been curated that easily expose a high Attack Success Rate (ASR) for both proprietary models [1] and publicly open-sourced models [31, 32]. On the other hand, without many bells and whistles, recent defense strategies - primarily focused on safety-aware supervised fine-tuning [71] and system prompt protection [60] - have surprisingly shown remarkable results on these benchmark datasets. In particular, VLLMs like LLaVA [31] can be fully safeguarded against the attacks involved (ASR $\rightarrow$ 0) [60, 71, 18]. As such, these dual findings raise an intriguing question: does it suggest that defending against jailbreak attacks is easy, given that the attacks themselves have already been shown to be relatively simple?
@@
The observation above highlights a perplexing safety paradox that undermines the credibility of recent advancements, particularly in defense mechanisms. To address this challenge, we present the first comprehensive study that underlines and elucidates the dual ease of both jailbreak attacks and defenses in VLLMs. Our first finding challenges prior assumptions that the vulnerability to jailbreak attacks stems from catastrophic forgetting or fine-tuning [71, 42]. Instead, we reveal that the true cause lies in the inclusion of image inputs, which compromises the guardrails of the base LLMs. On the other hand, we observe that existing defense mechanisms [71, 60] tend to be overly prudent. One typical manifestation is that VLLMs with post-defense, are prone to abstaining from responding even to benign queries. This issue of over-prudence significantly impairs the helpfulness of VLLMs, reducing their utility in real-world applications. In addition, we point out that the two well-studied evaluation methods often show a limited correlation in identifying jailbreaks, making it less convincing to determine the effectiveness of an attack method or a defense strategy.
@@
Beyond understanding the safety paradox, we further note that the jailbreak defense can be re-framed into a detection-then-response process. In view of this, we propose to implement an evaluator prior to the final VLLM response. We opt not to utilize an additional VLLM for detection as [18], given the limited reliability of current VLLMs in providing robust safeguards. Instead, we explore a vision-free evaluator, where we repurpose the guardrails of recent advanced LLMs to judge the harmfulness of a given textual query. Interestingly, we find that this evaluator, when paired with a VLLM for safe response generation, suffers less from the over-prudence problem, achieving a balanced interplay between robust safety alignment and model helpfulness.
@@
To the best of our knowledge, we are the first to investigate the safety paradox problem of VLLMs. Through the empirical findings presented in this work, we seek to highlight this issue and raise more attention to its significance. Beyond this, we hope to provide insights that can support future advancements in this field, such as reaching a consensus on the nature of jailbreak attacks and their associated risks, facilitating the collection of comprehensive attack data, and developing more robust evaluations.
2 Related Work
2.1 VLLM Attack
Existing jailbreak attacks on VLLMs can be broadly categorized into two groups: adversarial perturbation and prompt injection [30]. The former involves optimizing an adversarial image [36], either from random noise or a benign image, to elicit harmful responses [6, 49, 58, 46, 68]. The objective of this attack is to generate outputs that include a predefined list of toxic words. For instance, [19, 43] show that a single visual adversarial input can universally jailbreak an aligned VLLM. The imgJP [38] enables the attacks to be effective across a variety of unseen prompts and images. In contrast to these methods that operate within a constrained perturbation budget, prompt injection techniques deliberately manipulate image or instruction data without such limitations [9, 35, 61, 71]. The dominant techniques in this category focus on embedding high-risk content into images through typography or generative methods like stable diffusion [45]. For example, FigStep [17] utilizes textual prompts to induce MLLMs into completing sentences in an image that inadvertently result in malicious outputs step-by-step. MM-SafetyBench [33] generates harmful images spanning 13 commonly encountered scenarios. SASP [61] aims to hijack the system prompt by using GPT-4 [39] as a red teaming tool against itself, searching for potential jailbreak prompts.
2.2 VLLM Defense
2.3 LLM Attack and Defense
LLM jailbreak attack methods can be roughly classified into white-box and black-box attacks based on the transparency of the victim models [72, 62]. White-box attack strategies include efforts to search for jailbreak prompts by leveraging model gradients [25, 70, 72] or predicted logits of output tokens [67, 21]. Additionally, some methods involve fine-tuning the target LLMs with adversarial examples to induce harmful behaviors [27, 44, 66]. In contrast, prompt manipulation constitutes the primary method employed in the more challenging black-box attacks [65, 13, 63]. To defend against such jailbreak attacks, various approaches have been proposed, including safeguarding system prompts [48, 73], implementing supervised fine-tuning [4, 5, 12], and developing RLHF techniques [3, 40, 15, 54].
@@ 2.3 LLM Attack and Defense
Table 1: Statistics of the four evaluated jailbreak attack datasets. #HS: number of harmful scenarios, such as illegal activity; Quadrants correspond to those defined in Fig. 1.
@@
3 Preliminary
Evaluation methods. There are two key methods for evaluating the harmfulness of model outputs: rule-based and LLM-based approaches [62]. Rule-based methods assess the effectiveness of an attack by searching for specific keywords in the VLLMs’ responses [71, 60]. This approach hinges on the fact that rejection responses typically include phrases like ‘I’m sorry’, or ‘I cannot answer’. LLM-based methods, on the other hand, utilize a state-of-the-art LLM as the evaluator to determine the success of an attack [23]. In this approach, the prompt and the response generated by a jailbreak attack are input into the evaluator, which then provides either a binary judgment or a fine-grained score to represent the degree of harmfulness.
@@ 4 Ease of Jailbreak Attack
Figure 2: Examples of harmful captions generated by the QWen2-VL model [56] in response to benign, general prompts. Top: Hateful speech against specific religions; Bottom: Harmful racially biased history. More contentious cases, such as those involving sensitive political issues, are shown in the supplementary material.
Figure 3: T-SNE visualization of features from unsafe and safe instructions (the safe points are overlaid by unsafe ones for the 3rd, 6th, and 9th figure). Unlike the other two text-only models, VLLM-MM processes both textual instructions and images, where images are consistent for both safe and unsafe instructions. The safety alignment inherent in the original LLM-Base is maintained in VLLM-Text, but is significantly compromised in VLLM-MM.
(a) Image attention weight distribution from the last layer (left) and preceding layers (right).
(b) Image attention weights for safe (x-axis) and unsafe instructions (y-axis).
Figure 4: VLLMs pay more attention to unsafe images compared to safe images for benign instructions (a). However, when constrained to the same images, the distribution of attention weights remains almost the same (b) across instructions with distinct safety attributes.
Existing VLLMs have shown significant potential across a broad range of general tasks, including understanding, reasoning, and planning [50]. However, they are also notorious for their susceptibility to relatively simple attacks, particularly jailbreak attacks. To better understand the manifestation of this vulnerability and the underlying reasons, we analyze six state-of-the-art VLLMs (see Table 2) from several perspectives.
@@ 4.1 Evidence
Caption jailbreak. In addition to generating unsafe responses to harmful instructions, we observe that existing VLLMs can also produce inappropriate interpretations of images in response to benign, general caption prompts. For instance, we utilize a neutral caption prompt - Please describe the content of this image - which is not expected to elicit harmful or sensitive information, to query a VLLM. However, as shown in the two examples of Fig. 2, the model produces captions that spread hateful speech against certain religions and harmful racially biased history, respectively.
4.2 Rationale
Our explanation for the ease of jailbreak attacks on VLLMs, presented below, contrasts with the findings of previous studies [71, 42]:
@@ 4.2 Rationale
To explore this point, we conduct in-depth experiments on the VL-Guard dataset [71] using several VLLMs. The VL-Guard dataset provides two key advantages that support our findings: 1) Each harmless image is paired with both a harmful instruction and a harmless instruction. 2) The dataset maintains a balance between harmful and harmless images. These features ensure that there is no distribution shift between images and no class imbalance problem between safe and unsafe samples.
Our observations pertaining to jailbreak attacks are summarized into the following two points:
– VLLMs are unable to distinguish between safe and unsafe inputs, whereas their LLM base can. We visualize the encoded features of both safe and unsafe instructions from the final transformer layer in Fig. 3. For this experiment, we utilize three VLLMs, i.e., LLaVA-1.5-Vicuna-7B, LLaVA-NeXT-Mistral-7B, and LLaVA-NeXT-Llama3-8B, along with their corresponding LLM bases, Vicuna [69], Mistral [24], and Llama-3 [14]. It is important to note that the pre-trained weights from these base LLMs have been further fine-tuned by their respective VLLMs. The features are averaged across textual tokens for the LLM-Base and VLLM-Text, and across both textual and visual tokens for VLLM-MM.
@@
– VLLMs pay more attention to harmful images than harmless ones. We further investigate why VLLMs fail to abstain from following instructions for harmful images, even when it comes to simple captioning (Sec. 4.1). Fig. 4(a) shows the attention weights assigned to image tokens for benign instructions. It is evident that VLLMs tend to focus more on visual tokens from harmful images than from harmless ones, increasing the risk of generating unsafe content from these harmful images. We confirm that this effect is due to the harmfulness of the images themselves, rather than the harmfulness of the instructions. In detail, Fig. 4(b) demonstrates that when analyzing the same image, the attention weights for safe and unsafe instructions are nearly identical.
@@ 5 Ease of Jailbreak Defense
In contrast to the vulnerabilities of VLLMs in the face of jailbreak attacks, we arrive at another rather surprising and counterintuitive conclusion: VLLMs are, in fact, also relatively easy to defend against these very attacks. This insight is mainly motivated by recent studies that reveal how employing simple defense mechanisms can yield near-optimal performance on benchmark datasets [60, 71, 18]. The ease of these defenses, when juxtaposed with the apparent ease of attack, suggests a nuanced dynamic in the safety landscape of VLLMs. However, does this observation truly challenge the assumption that advanced models are inherently difficult to defend?
@@ 5.1 Evidence
We investigate two representative groups of methods in this experiment: safety-aware supervised fine-tuning Mixed VLGuard [71] and the training-free, prompt-based defense AdaShield-A [60]. The results of these methods are presented in Table 3 and Table 4, respectively (numbers are reproduced from the original papers). Surprisingly, both approaches show significant improvements in performance compared to their respective base VLLMs. Some VLLMs, such as LLaVA-1.5-13B on the FigStep benchmark in Table 3, achieve optimal safeguard. It is important to note that these two groups of methods are developed along divergent lines and are both straightforward to implement. Similar outcomes have also been observed in other defense studies [18]. These results suggest that current VLLMs are relatively easy to defend against jailbreak attacks, at least based on the superficial numbers observed across these benchmark datasets.
5.2 Rationale
We explain the observation that existing VLLMs are remarkably easy to defend against jailbreak attacks, with the following two remarks.
@@ 5.2 Rationale
Existing defense approaches [71, 60] demonstrate the effectiveness of further fine-tuning and system prompt protection on some limited datasets. However, it could be argued that these methods may not generalize to other jailbreak datasets. Our initial findings challenge this assumption, showing that these approaches extrapolate well to unseen datasets. Intrigued by these results, we then ask: how do they perform on benign inputs?
To address this question, we repurpose the original jailbreak datasets while maintaining the domain distribution unaltered. Specifically, for benign inputs lying in Quadrant I of Fig. 1, VLLMs are expected to respond without abstention [20]. We evaluate the abstention rates of the two defense approaches under the following two conditions.
@@
Our second point draws attention to the intrinsic limitations associated with the evaluation methods.
@@
Recall that the majority of evaluation methods consist of rule-based approaches (i.e., keyword matching) and model-based methods (e.g., Llama-Guard [23]). To quantify the level of agreement between these two approaches, we employ Cohen’s kappa statistic [11]. The upper bound of this value is 1, indicating perfect agreement between the two populations. Conversely, a value close to 0 or negative suggests that the methods share little to no consistency. The results are illustrated in Fig. 7. As can be observed in the figure, the values are predominantly negative or close to 0, indicating that the two methods fail to reach a consensus in most cases. Consequently, strong defense performance measured by one evaluation metric can be contradicted by results from the other.
5.3 A Simple Defense Baseline
In addition, we conduct experiments in which VLLMs are deliberately instructed to abstain from answering queries. Specifically, we require these models to always respond with ‘I’m sorry’ after answering questions. The experimental results on two datasets are presented in Fig. 8.
@@ 5.3 A Simple Defense Baseline
  • FigStep [17]. As shown in this figure, the explicit abstention prompt effectively ‘protects’ all three models. In particular, each model achieves an ASR approaching zero following this pseudo-defense strategy.
@@
  • MM-Vet [64] serves as a general multi-modal benchmark, distinct from FigStep by including only benign queries and images. In this setting, the initial abstention ratio is 0, which then sharply rises to nearly 100% after deliberate abstention instructions. Besides, we found that the instruction-following capability becomes a key factor in this context. Specifically, the recent, more robust model QWen2-VL achieves a 100% abstention rate, with a slight reduction in the original accuracy on MM-Vet. In contrast, the relatively inferior models, LLaVA-Next-Mistral and LLaVA-1.5-Vicuna, experience a modest performance decline.
@@
Figure 8: Variance in model performance after deliberately instructing VLLMs to abstain from answering. (Left) ASR results on FigStep, where lower values indicate better performance. (Right) Abstention ratio (lower values preferred) and accuracy on the general benchmark MM-Vet [64].
6 Vision-free Evaluator

### v2

@@ Abstract
The vulnerability of Vision Large Language Models (VLLMs) to jailbreak attacks appears as no surprise. However, recent defense mechanisms against these attacks have reached near-saturation performance on benchmark evaluations, often with minimal effort. This dual high performance in both attack and defense raises a fundamental and perplexing paradox. To gain a deep understanding of this issue and thus further help strengthen the trustworthiness of VLLMs, this paper makes three key contributions: i) One tentative explanation for VLLMs being prone to jailbreak attacks–inclusion of vision inputs, as well as its in-depth analysis. ii) The recognition of a largely ignored problem in existing defense mechanisms–over-prudence. The problem causes these defense methods to exhibit unintended abstention, even in the presence of benign inputs, thereby undermining their reliability in faithfully defending against attacks. iii) A simple safety-aware method–LLM-Pipeline. Our method repurposes the more advanced guardrails of LLMs on the shelf, serving as an effective alternative detector prior to VLLM response. Last but not least, we find that the two representative evaluation methods for jailbreak often exhibit chance agreement. This limitation makes it potentially misleading when evaluating attack strategies or defense mechanisms. We believe the findings from this paper offer useful insights to rethink the foundational development of VLLM safety with respect to benchmark datasets, defense strategies, and evaluation methods.
@@ 1 Introduction
Jailbreak attacks, also nowadays referred to as red-teaming [45], serve as the most common method for assessing the harmlessness of LLMs and Vision-LLMs (VLLMs) [57, 34, 8]. They are designed to circumvent the built-in restrictions or safeguards within models [25], eliciting them to produce malicious outputs, such as content related to illegal activities, hate speech, and pornography. Compared to their LLM counterparts, the vulnerability of VLLMs to jailbreak attacks has garnered considerable attention until very recently [65, 32]. Some initial methods inject high-risk content into images through typography or generative techniques like stable diffusion [49]. Leveraging such methods, datasets have been curated that easily expose a high Attack Success Rate (ASR) for both proprietary models [1] and publicly open-sourced models [33, 34].
@@
On the other hand, without many bells and whistles, recent defense strategies–primarily focused on safety-aware supervised fine-tuning [75] and system prompt protection [64]–have shown surprisingly remarkable defense results on these benchmark datasets. In particular, VLLMs like LLaVA1.5 [33] and MiniGPT-v2 [8] can be fully safeguarded against the attacks involved (ASR $\rightarrow$ 0) [64, 75, 20]. This dual-ease finding raises an intriguing question: does it suggest that defending against jailbreak attacks is easy, given that the attacks themselves have already been known to be relatively painless?
@@
The observation above highlights a perplexing safety paradox. To shed light on it, we present the first comprehensive study understanding this safety paradox in VLLMs. i) Our first finding challenges prior assumptions that the vulnerability to jailbreak attacks stems from catastrophic forgetting or fine-tuning [75, 46]. Instead, we reveal that the true cause lies in the inclusion of image inputs, which compromises the guardrails of the backbone LLMs. ii) On the other hand, we observe that existing defense mechanisms [75, 64] tend to be overly prudent. One typical manifestation is that VLLMs with post-defense, are prone to abstaining from responding even to benign queries. This issue of over-prudence significantly impairs the helpfulness of VLLMs. Even more concerning, we demonstrate that a simple, deliberate abstention approach–such as post-fixing a prompt Please respond “I’m sorry” after answering questions to each query–already yields favorable results for models with advanced instruction-following capabilities (i.e., InternVL-2 [11]). Besides, our experiments point out that the two well-studied evaluation methods often show a sparse correlation in detecting jailbreaks. Specifically, some attacks that are successfully identified by rule-based evaluations can often escape from LLM-based evaluations. This discrepancy weakens the accuracy of evaluating an attack method or a defense strategy.
@@
Beyond understanding the safety paradox, we note that the jailbreak defense can be re-framed into a detection-then-response process. iii) As such, we propose to implement a detector prior to the final VLLM response and design a simple plug-and-play LLM-Pipeline approach. We opt not to utilize an additional VLLM for detection as ECSO [20], given the limited reliability of current VLLMs in providing robust safeguards. Instead, we explore a vision-free detector, where we repurpose the guardrails of recent advanced LLMs (e.g., Llama3.1 [57]) to judge the harmfulness of a given textual query, optionally with the image caption. Interestingly, we find that this detector, when paired with a VLLM for safe response generation, suffers less from the over-prudence problem, achieving a balanced interplay between robust safety alignment and model helpfulness.
@@
To the best of our knowledge, we are the first to investigate the safety paradox problem of VLLMs. Through the empirical findings presented in this work, we seek to highlight this issue and raise more attention to its significance. Beyond this, we hope to provide insights that can support future advancements in this field, such as reaching a consensus on the nature of attacks and their associated risks, facilitating the collection of comprehensive attack data, and developing more robust defense and evaluations.
2 Preliminary
@@
Table 1: Statistics of the four evaluated jailbreak attack datasets. #HS: number of harmful scenarios, such as illegal activity and hate speech; Quadrants correspond to those defined in Fig. 1.
@@
Evaluation methods. There are two key methods for evaluating the harmfulness of model outputs: rule-based and LLM-based evaluations [66]. Rule-based methods assess the effectiveness of an attack by searching for specific keywords in the VLLMs’ responses [75, 64]. This approach hinges on the fact that rejection responses typically include phrases like ‘I’m sorry’, or ‘I cannot answer’. LLM-based methods, on the other hand, utilize a state-of-the-art LLM as the evaluator to determine the success of an attack [25]. In this approach, the prompt and the response generated by a jailbreak attack are input into the evaluator, which then provides either a binary judgment or a fine-grained score to represent the degree of harmfulness.
@@ 3 Ease of Jailbreak Attack
Figure 2: Examples of harmful captions generated by the QWen2-VL model [60] in response to benign caption prompts. Top: Hateful speech against specific religions; Bottom: Harmful racially biased history. More contentious cases, such as those involving sensitive political issues, are shown in the supplementary material.
Figure 3: T-SNE visualization of features from unsafe(U) and safe(S) instructions (the safe points are overlaid by unsafe ones for figures 3, 6, and 9). Unlike the other two text-only models, VLLM-MM processes both textual instructions and images. The safety alignment inherent in the original LLM-Base is maintained in VLLM-Text, but is significantly compromised in VLLM-MM.
(a) Image attention distribution from the last layer (left) and preceding layers (right).
(b) Image attention for safe (x-axis) and unsafe instructions (y-axis).
Figure 4: Image attention statistics from the [CLS] token of LLaVA. (a) For benign instructions, VLLMs pay more attention to unsafe images compared to safe images. (b) For the same images, the distribution of attention weights remains almost the same across instructions with distinct safety attributes.
Existing VLLMs have shown significant potential across a broad range of general tasks, including understanding, reasoning, and planning [54]. However, they are also notorious for their susceptibility to relatively simple attacks, particularly jailbreak attacks.
@@ 3.1 Evidence
Caption jailbreak. In addition to generating unsafe responses to harmful instructions, we observe that existing VLLMs can also produce inappropriate interpretations of images in response to benign, general caption prompts. For instance, we utilize a neutral caption prompt–Please describe the content of this image–which is not expected to elicit harmful or sensitive information, to query a VLLM. However, as shown in the two examples of Fig. 2, the model produces captions that spread hateful speech against certain religions and harmful racially biased history, respectively.
3.2 Rationale: Inclusion of Vision Inputs
Our explanation for the ease of jailbreak attacks on VLLMs contrasts with the findings of previous studies [75, 46]:
@@ 3.2 Rationale: Inclusion of Vision Inputs
To explore this point, we conduct in-depth experiments on the VL-Guard dataset [75] using several VLLMs. The VL-Guard dataset provides two key advantages that support our findings: 1) Each safe image is paired with both a harmful instruction and a safe instruction. 2) The dataset maintains a balance between harmful and safe images. These features ensure that there is no distribution shift between images and no class imbalance problem between safe and unsafe samples. Our observations are summarized into the following two points:
$\bullet$ VLLMs are unable to distinguish between safe and unsafe inputs, whereas their LLM base can.
We visualize the encoded features of both safe and unsafe instructions from the last transformer layer in Fig. 3. For this experiment, we utilize three VLLMs, i.e., LLaVA-1.5-Vicuna-7B, LLaVA-NeXT-Mistral-7B, and LLaVA-NeXT-Llama3-8B, along with their corresponding LLM bases, Vicuna [73], Mistral [26], and Llama-3 [16]. It is important to note that the pre-trained weights from these base LLMs have been further fine-tuned by their respective VLLMs. The features are averaged across textual tokens for the LLM-Base and VLLM-Text, and across both textual and visual tokens for VLLM-MM.
@@
$\bullet$ VLLMs attend more to harmful images than safe ones.
We further investigate why VLLMs fail to abstain from following instructions for harmful images, even when it comes to simple captioning (Sec. 3.1). Specifically, the results in Fig. 4(a) show the attention weights assigned to image tokens for benign instructions. It is evident that VLLMs tend to focus more on visual tokens from harmful images than from safe ones, increasing the risk of generating unsafe content from these harmful images. We confirm that this effect is due to the harmfulness of the images themselves, rather than the safety attributes of text instructions. In detail, Fig. 4(b) demonstrates that when analyzing the same image, the attention weights for safe and unsafe instructions are nearly identical.
@@ 4 Ease of Jailbreak Defense
Besides the above observation that VLLMs are highly vulnerable to jailbreak attacks, we arrive at a rather surprising and counterintuitive conclusion: VLLMs are, in fact, also relatively easy to defend against these very attacks. This insight is mainly motivated by recent studies that reveal how employing simple defense mechanisms can yield near-optimal performance on benchmark datasets [64, 75, 20]. The ease of these defenses, when juxtaposed with the apparent ease of attack, suggests a nuanced dynamic in the safety landscape of VLLMs.
@@ 4.1 Evidence
We investigate two representative groups of methods in this experiment: safety-aware supervised fine-tuning, e.g., Mixed VLGuard [75] and the training-free, prompt-based defense, e.g., AdaShield-A [64]. The results of these methods are presented in Table 3 and Table 4, respectively (numbers are reproduced from the original papers). Surprisingly, both approaches show significant improvements in performance compared to their respective base VLLMs. Some models, such as LLaVA-1.5-13B on the FigStep benchmark in Table 3, achieve optimal safeguard. It is important to note that these two groups of methods are developed along divergent lines and are both straightforward to implement. Similar outcomes have also been observed in other defense studies like ECSO [20]. These result indicate that, at least based on the numerical results observed across benchmark datasets, current VLLMs appear relatively easy to defend against jailbreak attacks.
4.2 Rationale 1: The Over-Prudence Problem
Our first explanation for the ease of jailbreak defense lies in the over-prudence problem:
@@ 4.2 Rationale 1: The Over-Prudence Problem
Existing defense approaches demonstrate the effectiveness on some limited datasets. However, it could be argued that these methods may not generalize to other jailbreak datasets. Our initial findings challenge this assumption, showing that these approaches extrapolate well to unseen datasets. Intrigued by these results, we then ask: how do they perform on benign inputs?
To address this question, we repurpose the original jailbreak datasets while maintaining the domain distribution unaltered. In particular, for benign inputs lying in Quadrant I of Fig. 1, VLLMs are expected to respond without abstention [22]. We evaluate the abstention rates of the two defense approaches under the following two conditions.
@@
4.3 Rationale 2: Evaluation Dilemma
Beyond the over-prudence problem, our second explanation reveals the intrinsic limitations associated with the evaluation methods:
@@ 4.3 Rationale 2: Evaluation Dilemma
Recall that the majority of evaluation methods consist of rule-based approaches (i.e., keyword matching) and model-based methods (e.g., Llama-Guard [25]). To quantify the level of agreement between these two approaches, we employ Cohen’s kappa statistic [13]. The upper bound of this value is 1, indicating perfect agreement between the two populations. Conversely, a value close to 0 or negative suggests that the methods share little to no consistency. As can be observed in Fig. 7, the values are predominantly negative or close to 0, indicating that the two methods fail to reach a consensus in most cases [44, 37]. Consequently, strong defense performance measured by one evaluation metric can be contradicted by results from the other.
A Simple Defense Baseline. Driven by this evaluation dilemma, we then investigate whether a simple system prompt protection can bypass the evaluation protocol, i.e., pretending to be a successful defense. To this end, we instruct VLLMs to deliberately abstain beyond answering queries, e.g., always respond with ‘I’m sorry’ after answering questions. The experimental results on two datasets are presented in Fig. 8.
@@
  • FigStep [19] is a typical jailbreak dataset. As shown in the figure, the explicit abstention prompt effectively ‘protects’ all three models. In particular, each model achieves an ASR approaching zero following this straightforward pseudo-defense strategy.
@@
  • MM-Vet [68] serves as a general multi-modal benchmark, distinct from FigStep by including only benign queries and images. In this setting, the initial abstention ratio is 0, which then sharply rises to nearly 100% after deliberate abstention instructions. Besides, we found that the instruction-following capability becomes a key factor in this context. Specifically, the recent, more robust model Intern-VL2 achieves a 100% abstention rate, with a slight reduction in the original accuracy on MM-Vet. In contrast, the relatively inferior models, LLaVA-Next-Mistral and LLaVA-1.5-Vicuna, experience a modest performance decline.
@@
Figure 8: Variance in model performance after deliberately instructing VLLMs to abstain beyond answering. (Left) ASR results on FigStep, where lower values indicate better performance. (Right) Abstention ratio (lower values preferred) and accuracy on the general benchmark MM-Vet [68].
5 Jailbreak Detector and Beyond
Vision-free Detector. We propose two variants of vision-free detectors:
  • Naive–Directly taking the original textual instruction only as input.
  • Scenario–Including a scenario-aware system protection prompt before the given instruction. For instance, the prompt can be If you feel like the question will induce potentially harmful information, about Illegal Activities, please reject to respond..
We explore four recent LLMs as candidate detectors: Mistral-7B [26], QWen2.5-14B [56], Llama3.1-8B, and Llama3.1-70B [16].
5.1 Detector Results on Jailbreak Datasets
The overall results for the four detectors are presented in Table 5 and Table 6 (values on the left side of symbol $\rightarrow$). From these tables, we draw the following two observations:
  • The vision-free detectors already achieve highly competitive results on jailbreak datasets. For instance, as shown in Table 5, Llama3.1-8B reaches an ASR of 0.7 on VLSafe, while Llama3.1-70B with the Scenario variant achieves an ASR of 2.8 on FigStep. Exceptions observed on VLGuard and MM-SafetyBench (Table 6) stem from instructions requiring joint image-text understanding.
  • The Scenario approach consistently outperforms its Naive counterpart by a significant performance margin in most cases. This finding suggests that informing LLMs explicitly about the potential for harmful scenarios enhances their confidence in identifying jailbreaks.
Caption Re-check. We note that queries from the other two jailbreak datasets, VLGuard [75] and MM-SafetyBench [36], demand a joint understanding of both image and instruction. To address the limitations of LLMs lacking access to visual information, we propose using QWen2-VL-7B [60] to generate captions for the provided images, enabling LLMs to utilize these captions as contexts.